### v2

# HATT: Hamiltonian Adaptive Ternary Tree for Optimizing Fermion-to-Qubit Mapping

Abstract
This paper introduces the Hamiltonian-Adaptive Ternary Tree (HATT) framework to compile optimized Fermion-to-qubit mapping for specific Fermionic Hamiltonians. In the simulation of Fermionic quantum systems, efficient Fermion-to-qubit mapping plays a critical role in transforming the Fermionic system into a qubit system. HATT utilizes ternary tree mapping and a bottom-up construction procedure to generate Hamiltonian aware Fermion-to-qubit mapping to reduce the Pauli weight of the qubit Hamiltonian, resulting in lower quantum simulation circuit overhead. Additionally, our optimizations retain the important vacuum state preservation property in our Fermion-to-qubit mapping and reduce the complexity of our algorithm from $O(N^{4})$ to $O(N^{3})$. Evaluations and simulations of various Fermionic systems demonstrate $5\sim 20\%$ reduction in Pauli weight, gate count, and circuit depth, alongside excellent scalability to larger systems. Experiments on the Ionq quantum computer also show the advantages of our approach in noise resistance in quantum simulations.
I Introduction
Simulating Fermionic systems, composed of particles such as electrons, protons, and neutrons [15], is an essential application area for quantum computing. These systems are integral to various critical physics models, including the electronic structure of molecules in quantum chemistry [26], the Fermi-Hubbard lattice model in condensed matter [1], neutrino evolution from astroparticle physics [3, 32, 7], etc. Being able to simulate large scale Fermionic systems could directly benefits material science, drug discovery, etc. However, Fermionic systems are intrinsically quantum and hard to simulate on classical computers on a large scale due to their exponential and super-exponential complexities. It can take tens of millions of cores in a supercomputer to perform ab initio electronic structure simulation [18], and it is getting harder to scale larger, limiting the scales scientists could study. Thus, using quantum computers to simulate Fermionic systems provides a crucial solution to the complexity barrier as quantum computers can simulate these systems with linear resource requirements [14].
To simulate a Fermionic system on a quantum computer requires an essential step, the Fermion-to-qubit mapping. This is because most quantum computers are built with qubits with different statistical properties than Fermions. A Fermion-to-qubit mapping bridges this gap by creating the Fermionic statistics in a qubit system. As depicted in Figure 1, a Fermionic Hamiltonian $\mathcal{H}_{\mathcal{F}}$ (on the left) is usually expressed by the creation and annihilation operators $\{a_{i}^{\dagger},a_{i}\}$, which do not naturally arise in a qubit system. A Fermion-to-qubit mapping will find Pauli strings to represent the creation and annihilation operators, converting the Fermionic Hamiltonian into a qubit Hamiltonian $\mathcal{H}_{\mathcal{Q}}$, which is a weighted sum of Pauli strings as shown in the middle of Figure 1. The qubit Hamiltonian can then be compiled into circuits to simulate the original Fermionic system [40, 24]. Note that Fermion-to-qubit mapping is unique, and different mappings will lead to significantly different circuit implementation costs even for the same input Fermionic system, as shown in Figure 1.
Figure 1: Overview of simulating Fermionic systems with quantum computers. The mapping on the upper half is better than that in the lower half because it generated a qubit Hamiltonian with fewer Pauli operators, and the circuit implementation cost is lower.
It is naturally desired for people to have Fermion-to-qubit mappings that can minimize the simulation circuit implementation cost. However, existing Fermion-to-qubit mappings are not yet satisfying and have significant further optimization potential. These mappings can be classified into two major types. The first type of approach is the constructive method, such as Jordan-Wigner transformation [22], Bravyi-Kitaev transformation [5], and ternary tree mapping [27, 20]. Although some can generate asymptotically optimal mapping, these methods do not consider the input Hamiltonian and always generate the Fermion-to-qubit mapping using the same operator pattern, leading to suboptimal actual mapping overhead. The second type of approach is the exhaustive search method, such as Fermihedral [25]. Fermihedral formulates the Fermion-to-qubit mapping search into a Boolean satisfiability (SAT) problem and uses an SAT solver to find the optimal mapping with minimal Pauli weight in the final qubit Hamiltonian. Although this approach can find better mappings by encoding the input Hamiltonian in the SAT instance, it has a scalability problem due to the SAT’s intrinsic hardness and fails to accommodate large-size practical cases.
In this paper, we aim to push the boundary of Fermion-to-qubit mapping forward and design a new compilation algorithm that can generate lower-cost mapping results while leveraging the input Hamiltonian information in a scalable manner. Formulating a good input-adaptive Fermion-to-qubit mapping is a complicated constrained optimization problem since valid Fermion-to-qubit mappings must satisfy certain constraints, and the input Hamiltonions can have different patterns across various Fermionic systems. Our key observation is that existing approaches do not fully exploit the potential of mathematical objects in this domain, including the variants in the data structures and the properties of Pauli algebra. In particular, we notice that the flexibility of the ternary tree data structure can be the key to our new algorithm. Indeed, ternary trees with different structures can yield valid but different Fermion-to-qubit mappings. It is possible to alter the ternary tree structure based on the input Fermionic Hamiltonian to yield low-cost Fermion-to-qubit mappings without exhaustive search.
To this end, we propose the Hamiltonian-Adaptive Ternary Tree (HATT) framework that can compile optimized Fermion-to-qubit mappings tailored to the input Fermionic Hamiltonian. First, HATT comes with a new bottom-up ternary tree generation algorithm that can incorporate the input Fermionic Hamiltonian information into the generated tree structure. Starting from the leaf nodes, our algorithm gradually adds parent nodes to construct the tree to maximize the cancellation during Pauli operator multiplication in the input Hamiltonian, thus minimizing the Pauli weight on each qubit. Second, we adapt our algorithm with a careful node selection and operator pairing process during tree construction. By traversing down and up the constructed tree, we can ensure that the resulting mapping will map the vacuum state in the Fermionic system to the all-zero state in the qubit mapping. This retains the important vacuum state preservation property for a Fermion-to-qubit mapping. Third, we further apply a caching strategy to the tree structure during construction to reduce the complexity of traversing up and down during operator pairing from worst $O(N)$ to $O(1)$ when mapping $N$-mode Fermionic systems. Our optimization decreases the total complexity of the algorithm from $O(N^{4})$ to $O(N^{3})$.
We evaluate HATT extensively against existing construction and exhaustive search methods on various Fermionic system models. The results show that HATT generates close-to-optimal mappings for small-size cases compared with the optimal mapping from Fermihedral [25] and even outperforms the approximately optimal solutions of Fermihedral. For larger-scale cases where Fermihedral fails, HATT outperforms those constructive methods (Jordan-Wigner [22], Bravyi-Kiteav [5], balanced ternary tree [20]) with $\sim 15\%$ reduction in Pauli weight, $5\sim 20\%$ reduction in circuit depth, and $10\sim 30\%$ reduction in $CNOT$ gate in the simulation quantum circuit. In addition, HATT achieves the lowest bias and variance in the noisy simulation experiments. On the real system study on IonQ Forte 1 quantum computer, HATT achieves the smallest variance and the second best average result (the best average result comes from Fermihedral’s optimal mapping).
The major contributions of this paper can be summarized as follows:
  1. 1.
    We propose HATT, a ternary tree-based Fermion-to-qubit mapping compilation framework to construct input-adaptive ternary tree Fermion-to-qubit mappings based on the Fermionic Hamiltonian in polynomial time.
  2. 2.
    We further improve our tree construction algorithm and successfully implement the vacuum state preservation, a desired property for Fermion-to-qubit mapping without increasing the search complexity or affecting the generated mapping performance.
  3. 3.
    By designing a map to cache the structure of the constructed tree, we reduce the overall complexity of our overall algorithm from $O(N^{4})$ to $O(N^{3})$ ($N$ is system size), achieving significant speed up for large cases.
  4. 4.
    Evaluation on several Fermionic systems shows HATT has much better scalability than exhaustive search and can outperform existing constructive methods on system-dependent performance, including circuit depth, gate count, and higher accuracy in noisy simulation and real-system experiments.
II Background
In this section, we introduce the essential background for understanding this paper, including Fermionic systems, Pauli string, Fermion-to-qubit mapping. More fundamental concepts of quantum mechanics and quantum computing can be found in [34, 31].
II-A Fermionic System
Generally speaking, a Fermionic system contains a set of Fermionic modes. Each $i^{th}$ Fermionic mode is captured by a $2$D Hilbert space called the Fock space $\mathcal{F}_{i}$, and a pair of creation and annihilation operators $(a_{i}^{\dagger},a_{i})$.
II-A1 Fock Space
According to the Pauli exclusion principle [33], a Fermionic mode has two exclusive states: unoccupied (|$\quantity(\ket{0}_{\mathcal{F}})$ or occupied by one Fermion (|$\quantity(\ket{1}_{\mathcal{F}})$. Its Fock space $\mathcal{F}$ is the $2$D Hilbert space $\mathrm{span}\quantity{\ket{0}_{\mathcal{F}},\ket{1}_{\mathcal{F}}}$, where {|$\{\ket{0}_{\mathcal{F}},\ket{1}_{\mathcal{F}}\}$ is called the Fock basis.
The Fock space of a $N$-mode Fermionic system is a $2^{N}$ Hilbert space generated by the tensor products of the Fock space of all $N$ Fermionic modes. The Fock basis is thus the tensor products of each basis:
|$\ket{e_{N-1},\dots,e_{0}}_{\mathcal{F}}=\bigotimes_{i=0}^{N-1}\ket{e_{i}}_{\mathcal{F}}\quad e_{i}=0\text{ or }1$
where |$\ket{e_{i}}_{\mathcal{F}}$ corresponds to the $i^{th}$ Fermionic mode.
II-A2 Creation and Annihilation Operator
Each Fermionic mode defines a pair of creation and annihilation operators $(a^{\dagger}_{i},a_{i})$. They act by increasing or decreasing the occupation number:
[…]
$a^{\dagger}_{i}\ket{1_{i}}_{\mathcal{F}}=a_{i}\ket{0_{i}}_{\mathcal{F}}=\mathbf{0}$
Notice that any annihilation operator $a_{i}$ that applies on the vacuum state |$\ket{0\dots 0}_{\mathcal{F}}$ produces $\mathbf{0}$: $a_{i}\ket{0\dots 0}_{\mathcal{F}}\equiv\mathbf{0}$.
II-A3 Fermionic Hamiltonian
Finally, the system’s Hamiltonian characterizes the time evolution behavior of a Fermionic system. With creation and annihilation operators, we can second quantize [13] the Hamiltonian of a Fermionic System by expressing it with the weighted sum of operator products. A simple 2-mode Fermionic system can have the following Hamiltonian:
$\mathcal{H}_{\mathcal{F}}=c_{0}a^{\dagger}_{0}a_{0}+c_{1}a^{\dagger}_{1}a_{1}+c_{2}a_{0}^{\dagger}a_{1}^{\dagger}a_{0}a_{1}$ (1)
II-B Pauli String
Pauli strings are the most basic concepts to describe a qubit system, as all $N$-length Pauli strings form an orthogonal basis for $N$-qubit Hermitians [31]. A Pauli string $S$ of a $N$-qubit system is defined as the tensor product of $N$ Pauli operators:
$S=\sigma_{N-1}\otimes\dots\otimes\sigma_{0}\quad\text{where }\sigma_{i}\in\{I,X,Y,Z\}$
We say it has a length $N$. Each Pauli operator $\sigma_{i}$ operates solely on the qubit $q_{i}$. $I$ is the identity operator, and $\{X,Y,Z\}$ are the Pauli matrices:
$X=\matrixquantity[0&1\\ 1&0],\ Y=i\matrixquantity[0&-1\\ 1&0],\ Z=\matrixquantity[1&0\\ 0&-1]$
II-B1 Format
A Pauli string could be written in two forms:
  • $N$-length string: the string lists operators from $\sigma_{N-1}$ to $\sigma_{0}$, such as $XYIZ=X\otimes Y\otimes I\otimes Z$
  • Compact form: identity operators ($I$) are omitted, and each non-identity Pauli operator is subscripted with the qubit it operates, such as $XYIZ=X_{3}Y_{2}Z_{0}$. Since operators are annotated, their order does not matter.
II-B2 Time Evolution Operator
The ultimate goal of quantum simulation on a digital quantum computer is to execute the time evolution operator $e^{-i\mathcal{H}_{\mathcal{Q}}t}$ of a given qubit Hamiltonian $\mathcal{H}_{\mathcal{Q}}$ with quantum circuits. It begins with decomposing $\mathcal{H}_{\mathcal{Q}}$ to a unique linear combination of $N$-length Pauli strings:
$\mathcal{H}_{\mathcal{Q}}=\sum c_{j}S_{j}\quad c_{j}\in\mathbb{R}$
The time evolution operator $e^{-i\mathcal{H_{Q}}t}$ is then approximated by Trotterization and number of time steps $n$:
[…]
① easily converts into a sequence of circuit snippets, each implementing a term $\exp(itc_{j}S_{j}/n)$ with basic gates. Figure 2 (a)$\sim$(e) shows how term $\exp(itc_{j}XYIZ/n)$ is converted: (a) Apply $\mathcal{Y}=\frac{1}{\sqrt{2}}\matrixquantity(1&i\\ i&1)$ gates to qubits that have $Y$ operator in the Pauli string ($q_{2}$), and apply $H$ to the ones have $X$ ($q_{3}$). This shifts the basis of the corresponding qubit from $Z$ into $Y$ or $X$. (b) Use $CNOT$ gates to entangle all qubits with non-identity Pauli operators ($q_{3},q_{2}$) to a target qubit ($q_{0}$). (c) Rotate the target qubit ($q_{0}$) with $R_{z}(2c_{j}t/n)$. (d)$\sim$(e) Apply inverse gates in (b) and (a).
Figure 2: Circuit snippet of operator $\exp(itc_{j}XYIZ/n)$. $q_{0}$ is selected as the target qubit.
II-B3 Pauli Weight
The Pauli weight of a Pauli string $w(S)$ is the number of non-identity operators in it. For example, the string $XYIZ$ has weight $3$. The Pauli weight of a qubit Hamiltonian $\mathcal{H}_{\mathcal{Q}}=\sum c_{i}S_{i}$ is defined as the sum of weights of all decomposed Pauli strings: $w(\mathcal{H}_{\mathcal{Q}})=\sum w(S_{i})$. As discussed in Section II-B2, the identity operators in a Pauli string do not generate any gates. Thus, the Pauli weight of a qubit Hamiltonian is approximately proportional to the number of gates in the circuit of the time evolution operator $e^{-i\mathcal{H_{Q}}t}$.
II-C Fermion-to-Qubit Mapping
To simulate a Fermionic system on a digital quantum computer, we have to bridge the Fermionic Hamiltonian $\mathcal{H}_{\mathcal{F}}$ to the qubit Hamiltonian $\mathcal{H}_{\mathcal{Q}}$. To do this, we use Fermion-to-qubit mapping to find a set of Pauli strings to represent the creation and annihilation operators while satisfying the Fermionic anticommutation relationship. The mapping first pairs the $N$ creation and $N$ annihilation operators into $2N$ Majorana operators $M_{j}$:
$a^{\dagger}_{j}=\frac{M_{2j}-iM_{2j+1}}{2}\quad a_{j}=\frac{M_{2j}+iM_{2j+1}}{2}$ (2)
Each Majorana operator $M_{j}$ maps to a $N$-length Pauli string $S_{j}$, thus mapping the creation/annihilation operators and the Fermionic Hamiltonian to the qubit Hamiltonian.
For example, using the well-known Jordan-Wigner (JW) Fermion-to-qubit mapping [22], the example Hamiltonian in Equation (1) is mapped to:
$\mathcal{H}_{\mathcal{F}}\mapsto\mathcal{H}_{\mathcal{Q}}$ $=\frac{2c_{0}+2c_{1}-c_{2}}{4}II+\frac{c_{2}-2c_{0}}{4}IZ$
$+\frac{c_{2}-2c_{1}}{4}ZI-\frac{c_{2}}{4}ZZ$
with
$a_{0}^{\dagger}\mapsto 0.5IX-0.5iIY$ $a_{0}\mapsto 0.5IX+0.5iIY$
$a_{1}^{\dagger}\mapsto 0.5XZ-0.5iYZ$ $a_{1}\mapsto 0.5XZ+0.5iYZ$
and the Majorana operators are:
$M_{0}=IX\quad M_{1}=IY\quad M_{2}=XZ\quad M_{3}=YZ$
III Hamiltonian Adaptive Ternary Tree
We aim to design an efficient and scalable compilation framework to generate low-overhead Fermion-to-qubit mappings tailored to the input Hamiltonians. Our approach exploits the properties of Pauli strings and a unique data structure, the ternary tree, to yield Pauli strings satisfying the constraints of a valid Fermion-to-qubit. The Pauli strings are generated to maximize the Pauli operator cancellation during the multiplication of Majorana operators, leading to low Pauli weight post-mapping qubit Hamiltonians.
III-A Extracting Majorana Operators from Ternary Tree
We first introduce the ternary tree and how to extract Majorana operators in the Fermion-to-qubit mapping from a ternary tree.
Figure 3: Example of a ternary tree and extracting Fermion-to-qubit mapping
III-A1 Ternary Tree
Ternary trees are a special type of tree. All nodes that are not leaves are denoted as internal nodes. Each internal node of a ternary tree has at most three child nodes. A ternary tree is complete if every internal node has exactly three children. If a complete ternary tree has $n$ internal nodes, it has $12n+$ 1 leaves. Each path from the root to a leaf is always unique since a ternary tree does not have cycles. For example, the four red nodes in Figure 3 (a) are the internal nodes of a ternary tree, and Figure 3 (b) shows a complete ternary tree after the leaves (blue nodes) are added to ensure each internal node has three child nodes.
III-A2 Extract Pauli Strings from Ternary Tree
A ternary tree Fermion-to-qubit mapping for an $N$-mode Fermionic system can be constructed via $2N$ Pauli strings representing the $2N$ Majorana operators. The $2N$ Pauli strings are extracted from a complete ternary tree with $N$ internal nodes and $12N+$ 1 leaves. Each Pauli string is represented by a path from the root node to one leaf node in the ternary tree. This process is illustrated using the example in Figure 3.
We denote each internal node by $In_{j}$ where $j=0,\dots,N-1$. The internal node $In_{j}$ corresponds to the qubit $q_{j}$ in the qubit system to be mapped onto. The leaf nodes are denoted by $\{L_{i}\}$. In this example, the ternary tree has four internal nodes $In_{0}\sim In_{3}$ and nine leaves $L_{0}\sim L_{8}$.
Now, we take a path from the root node to one leaf. At each step in this path, there are three possible branches to select the next node because each internal node in a complete ternary tree has three children. The three branches are labeled by $X$, $Y$, and $Z$. Then, a Pauli string can be determined by following the path. Suppose the Pauli string $S_{i}$ is constructed by the path from the root to the leaf node $L_{i}$, and the path is:
$Path=In_{a}\text{(root)}\rightarrow In_{b}\rightarrow In_{c}\rightarrow\dots\rightarrow L_{i}$
Each internal node $In_{j}$ contributes a Pauli operator in the Pauli string $S_{i}$. For the Pauli operator on qubit $q_{j}$, If $In_{j}\notin Path$, the operator is $I_{j}$. If $In_{j}\in Path$, based on the branch $In_{j}\rightarrow In_{k}$ it takes, the node contributes $X_{j}$, $Y_{j}$, or $Z_{j}$ if $In_{k}$ is the left, middle, or right child of $In_{j}$. $S_{i}$ equals the tensor product of all these operators to which the nodes contribute. For example, the green nodes in Figure 3 (c) represent a path from the root node to one leaf in the complete ternary tree. The path starts from node $In_{2}$ to node $In_{0}$ via the branch $Y$, so we have $Y_{2}$ in the Pauli string. Then the path goes from node $In_{0}$ to node $In_{1}$ via the branch $Z$, so we have $Z_{0}$. Similarly, we have $X_{1}$ because of the $X$ branch from $In_{0}$ to the leaf, and $I_{3}$ because $In_{3}$ is not in this path. This Pauli string is $I_{3}Y_{2}X_{1}Z_{0}$.
As there are $12N+$ 1 leaves, paths from the root to these leaves can construct $12N+$ 1 distinct Pauli strings. It can be proved that any $2N$ Pauli strings out of the $12N+$ 1 strings satisfy all the constraints required for the $2N$ Majorana operators in a valid Fermion-to-qubit mapping [27, 20]. Thus, we can always construct a Fermion-to-qubit mapping for $N$ Fermionic modes through a complete ternary tree with $N$ internal nodes and $12N+$ 1 leaves.
Figure 4: $3$-mode ternary tree Fermion-to-qubit examples. (a) Balanced ternary tree. (b) Unbalanced ternary tree.
III-B Motivation: Drawback of Balanced Ternary Tree
Since the Pauli strings are determined by the paths from the root to the leaves in a ternary tree, the tree’s structure will determine the generated Pauli strings and affect the overhead when simulating the corresponding qubit Hamiltonian. The number of non-identity Pauli operators in one Pauli string is the number of internal nodes in this path. Therefore, to reduce the Pauli weight of the Pauli strings, previous work[20] employs the balanced ternary tree, which has the lowest depth on average. For $N$ Fermonic modes, it generates Pauli strings with an average Pauli weight of $\lceil\log_{3}(2N+1)\rceil$, which can be proved to be the theoretically optimal Pauli weight per Pauli string.
However, in this paper, we make a key observation that the minimal Pauli weight in the Pauli strings may not lead to the minimal circuit implementation overhead in the qubit Hamiltonian simulation. An actual Hamiltonian can be a complex expression of the Pauli strings. The Pauli operators may cancel with each other and become the identity during the multiplication of Pauli strings. A motivation example is shown in Figure 4.
Suppose we have a toy model $3$-mode Fermionic Hamiltonian (in Majorana operators) $\mathcal{H}_{\mathcal{F}}=c_{1}M_{0}M_{5}+c_{2}M_{1}M_{3}$. the mapping generated from a balanced ternary tree, as shown in Figure 4 (a), maps it to $\mathcal{H}_{\mathcal{Q}}=c_{1}(X_{0}X_{1})(Y_{0}Z_{2})+c_{2}(X_{0}Y_{1})(Y_{0}X_{2})=c_{1}^{\prime}Z_{0}X_{1}Z_{2}+c_{2}^{\prime}Z_{0}Y_{1}X_{2}$. The Pauli weight is 6. But the unbalanced ternary tree mapping, as shown in Figure 4 (b), can yield a lower Pauli weight. It maps the Hamiltonian to $\mathcal{H}_{\mathcal{Q}}=c_{1}(X_{0})(Y_{0}Z_{1})+c_{2}(Y_{0}X_{1}X_{2})(Y_{0}X_{1}Z_{2})=c_{1}^{\prime\prime}Z_{0}Z_{1}+c_{2}^{\prime\prime}Y_{2}$, which has a Pauli weight of 3.
The example above shows that a mapping generated by unbalanced ternary trees can have larger Pauli weights in the Pauli strings and smaller Pauli weights in the final qubit Hamiltonian because of the multiplication of the Pauli operators. This motivates us to construct the ternary tree adaptively based on the input Fermionic Hamiltonian, which can further decrease the Pauli weight in the final qubit Hamiltonian.
III-C Hamiltonian Adaptive Ternary Tree Construction
This section introduces our ternary tree construction algorithm that can exploit cancellation between Majorana operators for an input Fermionic Hamiltonian. Suppose we have a $N$-mode Fermionic Hamiltonian. Our compilation takes a bottom-up approach, starting with all the $12N+$ 1 leaves. We gradually append parent nodes to the tree and finally reach the root. These parent nodes are always the internal nodes representing qubits. Our algorithm will search for a tree structure to lower the Pauli weight on each qubit. The pseudo-code is in Algorithm 1, and an algorithm overview is in Figure 5. The algorithm details are explained below.
Algorithm 1 Hamiltonian Adaptive Ternary Tree Construction
$\mathcal{H}_{\mathcal{F}}$: Hamiltonian of the Fermionic system $\{S_{i}\}$: $2N$anticommute Pauli strings generated by the ternary tree, used as $2N$Majorana operators.
1:$\mathcal{H}_{\mathcal{Q}}\leftarrow\mathtt{preprocess}(\mathcal{H}_{\mathcal{F}})$
[…]
3:$w\leftarrow\infty$
[…]
5:$w^{\prime}\leftarrow\mathtt{pauli\_weight}(\mathcal{H}_{\mathcal{Q}},(O_{X},O_{Y},O_{Z}),i)$
[…]
7:$w\leftarrow w^{\prime}$
[…]
9:$\mathcal{U}.\mathtt{remove}(O_{X})$; $\mathcal{U}.\mathtt{remove}(O_{Y})$; $\mathcal{U}.\mathtt{remove}(O_{Z})$
10:$\mathcal{U}.\mathtt{add}(O_{2N+1+i})$
11:$O_{2N+1+i}.(X,Y,Z)\leftarrow(O_{X},O_{Y},O_{Z})$
12:$\mathcal{H}_{\mathcal{Q}}\leftarrow\mathtt{reduce}(\mathcal{H}_{\mathcal{Q}},(O_{X},O_{Y},O_{Z})\mapsto O_{2N+1+i})$
13$\triangleright$ reduce Hamiltonian $\mathcal{H}_{\mathcal{Q}}$EndFor
14:$O_{root}\leftarrow\mathcal{U}.\mathtt{pop}()$
15:$\{S_{i}\}\leftarrow\mathtt{walk\_tree}(O_{root})$
16:return $\{S_{i}\}$
[…]
Figure 5: Algorithm for ternary tree construction
III-C1 Setup
The algorithm compiles a complete ternary tree Fermion-to-qubit mapping for an $N$-mode input Fermionic Hamiltonian with $12N+$ 1 leaf nodes and $N$ internal nodes. For simplicity, all the nodes are denoted by $O_{i}$ in the rest of this paper. The $12N+$ 1 leaf nodes are denoted by $O_{0}\sim O_{2N}$. The $N$ internal nodes are denoted by $O_{2N+1}\sim O_{3N}$. Leaf node $O_{0\leq\mathbf{i}\leq 2N}$ (previously $L_{\mathbf{i}}$ in the ternary tree example in Figure 3) corresponds to the Pauli string $S_{\mathbf{i}}$ and is used as the Majorana operator $M_{i}$ in the Fermion-to-qubit mapping. Internal node $O_{2N+1\leq 2N+1+\mathbf{i}\leq 3N}$ (previously $In_{\mathbf{i}}$ in Figure 3) corresponds to qubit $q_{\mathbf{i}}$. The algorithm starts with $12N+$ 1 leaves in the ternary tree as the initial node set $\mathcal{U}=\{O_{0},\dots,O_{2N}\}$.
The initial Fermionic Hamiltonian $\mathcal{H_{F}}$ is $\mathtt{preprocessed}$ to substitute all creation and annihilation operators with Majorana operators $\{M_{i}\}$, then $\mathcal{H_{F}}$ is mapped to a qubit Hamiltonian $\mathcal{H_{Q}}$ by using Pauli strings to represent Majorana operators: $M_{i}\rightarrow S_{i}$. The algorithm finds a concrete Pauli string for each $S_{i}$. For example, we have a 3-mode Hamiltonian:
$\mathcal{H}_{\mathcal{F}}=$ $\ a_{0}^{\dagger}a_{0}+2a_{1}^{\dagger}a_{2}^{\dagger}a_{1}a_{2}$ (3)
$=$ $\ 0.5i\cdot M_{0}M_{1}-0.5i\cdot M_{2}M_{3}-0.5i\cdot M_{4}M_{5}$
$+0.5\cdot M_{2}M_{3}M_{4}M_{5}$
The algorithm begins with $\mathcal{U}=\{O_{0},\dots,O_{6}\}$, each carries $S_{0},\dots,S_{6}$.The Hamiltonian is preprocessed to $\mathcal{H_{Q}}=0.5i\cdot S_{0}S_{1}-0.5i\cdot S_{2}S_{3}-0.5i\cdot S_{4}S_{5}+0.5\cdot S_{2}S_{3}S_{4}S_{5}$.
III-C2 Iteration
The algorithm takes $N$ iterations, from 00 to $N-1$. In the $i^{th}$ step, the algorithm picks three nodes from $\mathcal{U}$ and grows a parent $O_{2N+1+i}$ for them. The three nodes are removed from $\mathcal{U}$, and $O_{2N+1+i}$ is added to $\mathcal{U}$, which reduces the size of $\mathcal{U}$ by 2. This settles the operator on qubit $q_{i}$ for all strings. The three nodes are carefully selected to produce a minimum Hamiltonian Pauli weight on $q_{i}$. Details are discussed as follows:
  • Node Selection (line $6\sim 12$ in Algorithm 1): Suppose we select three nodes $O_{X}$, $O_{Y}$, $O_{Z}$ as the $X,Y,Z$ child for node $O_{2N+1+i}$. Based on definition, Pauli strings $S_{X},S_{Y},S_{Z}$ correspond to nodes $O_{X},O_{Y},O_{Z}$ have $X,Y,Z$ operators on qubit $q_{i}$ while the rest always have $I$, as shown in Figure 6
    Figure 6: Picked nodes $O_{X},O_{Y},O_{Z}$, their parent $O_{2N+1+i}$ and corresponding Pauli operators on qubit $q_{i}$
    For each possible selection of $O_{X},O_{Y},O_{Z}$, we calculate the Pauli weight for the Hamiltonian on qubit $q_{i}$ ($\mathtt{pauli\_weight}$ in Algorithm 1 line 7). It can be done by only calculating the $i^{th}$ Pauli operator of the Hamiltonian. Based on the Pauli weight and greedy strategy, $O_{X},O_{Y},O_{Z}$ should give the minimum Pauli weight on qubit $q_{i}$ among all selections.
    For the example Hamiltonian in Equation (3), we will pick $O_{0},O_{1},O_{6}$ in the first step. Only $S_{0},S_{1},S_{6}$ will have $X,Y,Z$ operator on qubit $q_{0}$ while the rest have $I$, as shown in Figure 7 . The Pauli weight is:
    $S_{0}S_{1}\mapsto XY=Z(1)$
    $S_{2}S_{3},\ S_{4}S_{5}\mapsto II=I(0)$
    $S_{2}S_{3}S_{4}S_{5}\mapsto II=I(0)$
[…]
  • Update tree and node set $\mathcal{U}$ (line $13\sim 15$ in Algorithm 1): Remove $O_{X},O_{Y},O_{Z}$ from $\mathcal{U}$ and add $O_{2N+1+i}$ to $\mathcal{U}$. Connect $O_{X},O_{Y},O_{Z}$ to the $X,Y,Z$ child of node $O_{2N+1+i}$. The node set $\mathcal{U}$ for the example is updated as shown in Figure 7 .
    Figure 7: Example: select $O_{0},O_{1},O_{6}$ nodes and construct their parent node $O_{7}(q_{0})$. Thus, $S_{0},S_{1}$ and $S_{6}$ has $X_{0},Y_{0}$ and $Z_{0}$ in it. All the rest of Pauli strings get $I_{0}$; update node set $\mathcal{U}$ by removing $O_{0},O_{1},O_{6}$ and inserting $O_{7}$; reduce Hamiltonian $\mathcal{H}_{\mathcal{Q}}$ by the relationship: $O_{7}$ carries $S_{7}$, which is the common part of $S_{0},S_{1}$ and $S_{6}$, by $\{S_{0},S_{1},S_{6}\}=S_{7}\otimes\{X,Y,Z\}$. All other nodes carried another Pauli string $S_{i}^{\prime}$ such that $S_{i}=S_{i}^{\prime}\otimes I_{0}$.
  • Reduce Hamiltonian ($\mathtt{reduce}$ in Algorithm 1 line 16): The node $O_{2N+1+i}$ carries a Pauli string $S_{2N+1+i}$ such that:
    $S_{\{X,Y,Z\}}=S_{2N+1+i}\otimes\{X,Y,Z\}_{i}$
    For the rest of the nodes $\forall O_{j}\notin\{O_{X},O_{Y},O_{Z}\}$, we can also find a Pauli string $S_{j}^{\prime}$ such that $S_{j}=S_{j}^{\prime}\otimes I_{i}$. We update $O_{j}$ to carry $S_{j}^{\prime}$ instead of $S_{j}$.
    Then, we subsititue $S_{j}$ with $S_{j}^{\prime}$ and $\{S_{X},S_{Y},S_{Z}\}$ with $S_{2N+1+i}$ in $\mathcal{H_{Q}}$, and remove operators on qubit $q_{i}$ from $\mathcal{H}_{\mathcal{Q}}$ since it is settled for all Pauli strings and will not affect total Pauli weight.
    In the example, we will introduce node $O_{7}$ representing qubit $q_{0}$ that carries $S_{7}$. The relationships between $S_{7}$ and $S_{0},S_{1},S_{6}$ and between $S_{i}^{\prime}$ and $S_{i}$ are also shown in Figure 7 . By substituting using the relationship:
    $S_{0}\mapsto S_{7}\otimes X_{0}\quad S_{1}\mapsto S_{7}\otimes Y_{0}\quad S_{6}\mapsto S_{7}\otimes Z_{0}$
    $S_{i}\mapsto S_{i}^{\prime}\otimes I_{0}$
    each term of Hamiltonian $\mathcal{H}_{\mathcal{Q}}$ in Equation (3) reduces to:
    $S_{0}S_{1}\mapsto S_{7}S_{7}=I\otimes Z_{0}$
    $S_{2}S_{3}\mapsto S_{2}^{\prime}S_{3}^{\prime}\otimes I_{0},\ S_{4}S_{5}\mapsto S_{4}^{\prime}S_{5}^{\prime}\otimes I_{0}$
    $S_{2}S_{3}S_{4}S_{5}\mapsto S_{2}^{\prime}S_{3}^{\prime}S_{4}^{\prime}S_{5}^{\prime}\otimes I_{0}$
    Operator on $q_{0}$ ($\dots\otimes\sigma_{0}$ part) is settled and can be removed since it does not further affect the total Pauli weight. $\mathcal{H}_{\mathcal{Q}}$ reduces to:
    $\mathcal{H}_{\mathcal{Q}}^{\prime}=0.5i\cdot S_{2}^{\prime}S_{3}^{\prime}-0.5i\cdot S_{4}^{\prime}S_{5}^{\prime}+0.5\cdot S_{2}^{\prime}S_{3}^{\prime}S_{4}^{\prime}S_{5}^{\prime}$
    In the next iteration step, we have node set $\mathcal{U}=\{O_{2},O_{3},O_{4},O_{5},O_{7}\}$ and the reduced Hamiltonian $\mathcal{H}_{\mathcal{Q}}^{\prime}$. We can then repeat the above steps with the node set $\mathcal{U}$ and $\mathcal{H}_{\mathcal{Q}}^{\prime}$ until all $N=3$ iterations are performed.
III-C3 Termination
The algorithm always terminates after $N$ iterations since the initial size of $\mathcal{U}$ is $12N+$ 1 and reduces by two in each step. The final only node in $\mathcal{U}$ is the root of the ternary tree.
III-C4 Complexity
We disregard the complexity of calculating the Pauli weight in each iteration step since it is input-dependent. We estimate its overhead to be a constant determined by the input Hamiltonian. In the algorithm, we compared all possible selections of $O_{X},O_{Y},O_{Z}$, which are $\binom{N}{3}\sim N^{3}$ choices. Thus, the complexity of each step is $O(N^{3})$. For total $N$ steps, the complexity is $O(N^{4})$.
IV Vacuum State Preservation and Performance Optimization
The algorithm described in Section III-C decreases the Hamiltonian Pauli weight but fails to retain one desired property of a Fermion-to-qubit mapping: vacuum state preservation. In addition, the computation complexity $O(N^{4})$ is still high. In this section, we further optimize our algorithm to a) ensure vacuum state preservation in the generated mapping and b) reduce its complexity to $O(N^{3})$.
IV-A Vacuum State Preservation
In Fermion-to-qubit mappings, it is desired to map the vacuum state of a Fermionic system |$\ket{0,\dots,0}_{\mathcal{F}}$ to the zero qubit state |$\ket{0}^{\otimes N}$ of the qubits. It is called Vacuum State Preservation, allowing lower overhead for state preparation in quantum simulation. This property is achieved by ensuring that the annihilation operators always produce $\mathbf{0}$ when applying on the vacuum state:
$\forall j,a_{j}\ket{0,\dots,0}_{\mathcal{F}}=\mathbf{0}\Leftrightarrow\frac{M_{2j}+iM_{2j+1}}{2}\ket{0}^{\otimes N}=\mathbf{0}$
To satisfy the right-hand-side equation, the corresponding Pauli strings $S_{2j},S_{2j+1}$ of Majorana operators $M_{2j},M_{2j+1}$ should have a $(X,Y)$ Pauli operator pair on a qubit, since $((X+iY)/2)\ket{0}\equiv\mathbf{0}$, and all the rest operators act the same on |$\ket{0}$, i.e., the $k^{th}$ Pauli operator $\sigma_{k}^{2j},\sigma_{k}^{2j+1}$ of $S_{2j},S_{2j+1}$ satisfies $\sigma_{k}^{2j}\ket{0}=\sigma_{k}^{2j+1}\ket{0}$ [27]. In the vanilla balance ternary tree [27, 20], this is achieved by re-assigning the Pauli strings to Majorana operators.
IV-B Operator Pairing In Tree Construction
However, we cannot re-assign Pauli strings to Majorana operators in our Hamiltonian-adaptive ternary tree construction because we have already assumed the Pauli string $S_{j}$ is assigned to the Majorana operator $M_{j}$. Reassignment destroys the Pauli operator cancellation created in our algorithm.
Instead of re-assigning Pauli strings, we improve our tree construction algorithm by enforcing vacuum state preservation during the node selection. That is, we only freely select $O_{X}$ and $O_{Z}$, and $O_{Y}$ is determined based on $O_{X}$ and $O_{Z}$. With careful construction leveraging the property of ternary trees, we can guarantee that all the Majorana operator pairs $(M_{2j},M_{2j+1})$ have an $(X,Y)$ Pauli operator pair on one qubit, and operators on other qubits act the same on |$\ket{0}$, and thus ensure vacuum state preservation.
We first define two new concepts in our new algorithm:
  1. I)
    $Z$-descendant: The $Z$-descendant of a node $O$, denoted as $desc_{Z}(O)$, is defined as the rightmost leaf of the subtree with root $O$. It can be reached by traversing down all the $Z$ branches. If $O$ is a leaf, then $desc_{Z}(O)=O$.
  2. II)
    Valid Pair: Two Pauli strings $(S_{2j},S_{2j+1})$ forms a valid pair if they share a $(X,Y)$ pair on a qubit and all other operators act the same on |$\ket{0}$. If we have $(S_{2j},S_{2j+1})$ is a valid pair forall $0\leq j<N$, then vacuum state preservation is guranteed.
We now introduce our improved tree construction algorithm 2. The overall structure is similar to Algorithm 1, but node selection is changed. In each step $i$, as shown in Algorithm 2 instead of selecting all $O_{X},O_{Y},O_{Z}$ nodes, we now only select nodes $O_{X},O_{Z}$ as the $X,Z$ children of node $O_{2N+1+i}$ and find an appropriate $O_{Y}$ to ensure exists $l$, such that $desc_{Z}(O_{X})$ corresponds to $S_{2l}$ and $desc_{Z}(O_{Y})$ corresponds to $S_{2l+1}$. Thus, $(S_{2l},S_{2l+1})$ is valid paired since they have a $(X,Y)$ pair on qubit $q_{i}$, and for all $k\neq i$, the $k^{th}$ operators for $S_{2l},S_{2l+1}$ are either the same or a $Z,I$ pair, whereas $Z\ket{0}\equiv I\ket{0}$, as shown in Figure 8.
Algorithm 2 Optimized HATT
$\mathcal{H}_{\mathcal{F}}$: Hamiltonian of the Fermionic system $\{S_{i}\}$: $2N$Majorana operators generated by the ternary tree
[…]
5:$O_{y}\leftarrow O_{x+1}$
6:$O_{Y}\leftarrow\mathtt{traverse\_up}(O_{y},\mathcal{U})$Else
7:$O_{y}\leftarrow O_{x-1}$
8:$O_{Y}\leftarrow\mathtt{traverse\_up}(O_{y},\mathcal{U})$
9:$\mathtt{swap}(O_{X},O_{Y})$triangleright swap $Y,X$ to $X,Y$EndIf
10$\cdots$ $\triangleright$ Algorithm 1, line $7\sim 11$EndFor
11$\cdots$ $\triangleright$ Algorithm 1, line $13\sim 16$EndFor
12$\cdots$ $\triangleright$ Algorithm 1, line $18\sim 20$
13:
[…]
1:$O\leftarrow O.Z$EndWhile
2:return $O$EndProcedure
3:
[…]
$\mathtt{traverse\_up}$$O,\mathcal{U}$$O\notin\mathcal{U}$
1:$O\leftarrow O.\text{parent}$EndWhile
2:return $O$EndProcedure
[…]
Suppose $O_{X}$ and $O_{Z}$ are selected. We traverse down along the $Z$ child from node $O_{X}$ until reaching a leaf of the final leaf $O_{x}=desc_{Z}(O_{X})$ (where $10\leq x<2N+$ 1), as shown in Figure 8 ①. Then, we select its nearby leaf $O_{y}$ as $desc_{Z}(O_{Y})$ based on $O_{x}$ to ensure the corresponding Pauli strings $S_{x},S_{y}$ form a valid pair $(S_{x=2l},S_{y=2l+1})$, as shown in Figure 8 ②:
  • If $x=2N$, then $O_{x}$ is the rightmost leaf. Discard this selection and pick $O_{X},O_{Z}$ again ($S_{2N}$ is always discarded in the ternary tree mapping and does not pair).
  • If $x$ is even ($x=2l$), let $y=x+1$ ($y=2l+1$).
  • If $x$ is odd ($x=2l+1$), let $y=x-1$ ($y=2l$). We must also exchange $O_{X}$ with $O_{Y}$ when we get $O_{Y}$ (Line 15, Algorithm 2) to ensure $(S_{x},S_{y})$ has a $(X,Y)$ Pauli operator pair, instead of $(Y,X)$.
To find $O_{Y}$ based on $O_{y}=desc_{Z}(O_{Y})$, we can traverse up from $O_{y}$ until we reach a node $O_{Y}$ in the current node set $\mathcal{U}$, as shown in Figure 8 ③. $O_{Y}$ is the ancestor of $O_{y}$, as it is the root of the subtree in which $O_{y}$ is located.
Figure 8: Procedure of finding $O_{Y}$ based on $O_{X},O_{Z}$
Finally, $O_{Y}$ is chosen as the $Y$ child of $O_{2N+1+i}$ to ensure valid string pairing. For each selection of $O_{X},O_{Z}$ in step $i$, we calculate the Pauli weight similarly to the original algorithm and construct the node based on the selection of $O_{X},O_{Y},O_{Z}$ that gives a minimum Pauli weight on qubit $q_{i}$, then reduce the Hamiltonian similarly to the original algorithm.
Consider our previous example, where $\mathcal{H}_{\mathcal{F}}=a_{0}^{\dagger}a_{0}+2a_{1}^{\dagger}a_{2}^{\dagger}a_{1}a_{2}=0.5i\cdot M_{0}M_{1}-0.5i\cdot M_{2}M_{3}-0.5i\cdot M_{4}M_{5}+0.5\cdot M_{2}M_{3}M_{4}M_{5}$. We have already selected $O_{0},O_{1},O_{6}$ as the children of $O_{7}$ in the first step, and the Hamiltonian is reduced to $\mathcal{H_{Q}}=0.5iS_{2}^{\prime}S_{3}^{\prime}-0.5iS_{4}^{\prime}S_{5}^{\prime}+0.5S_{2}^{\prime}S_{3}^{\prime}S_{4}^{\prime}S_{5}^{\prime}$.
In the second step, we first choose $O_{7},O_{2}$ as $O_{X},O_{Z}$, as shown in Figure 9 (a). However, $O_{x}=desc_{Z}(O_{7})=O_{6}$ is the rightmost operator, so we discard this selection and move on. Then, we choose $O_{2},O_{7}$ as $O_{X},O_{Z}$, as shown in Figure 9 (b). Here, $O_{x}=desc_{Z}(O_{2})=O_{2}$, thus $O_{y}=O_{2+1}=O_{3}$ and traverse back to $S$ gives $O_{Y}=O_{3}$. This choice also minimizes the Pauli weight. $O_{2},O_{3},O_{7}$ are the children of $O_{8}$.
We then follow the original procedure and reduce the Hamiltonian:
$S^{\prime}_{2}S^{\prime}_{3}\mapsto XY\rightarrow Z(1)$
$S^{\prime}_{4}S^{\prime}_{5}\mapsto II\rightarrow I(0)$
$S^{\prime}_{2}S^{\prime}_{3}S^{\prime}_{4}S^{\prime}_{5}\mapsto XYII\rightarrow Z(1)$
[…]
We can check that Majorana operators are paired validally. $(M_{2},M_{3})\mapsto(S_{2},S_{3})$ have a $(X,Y)$ pair on qubit $1$, and $(M_{0},M_{1})\mapsto(S_{0},S_{1})$ have a $(X,Y)$ pair on qubit $q_{0}$.
Figure 9: An example of operator pairing in tree construction. In (a), we select $O_{7},O_{2}$ as $O_{X},O_{Z}$, but $O_{x}=desc_{Z}(O_{7})=O_{6}$ is the rightmost leaf, thus we discard this selection. In (b), we select $O_{2},O_{7}$ as $O_{X},O_{Z}$ and found $O_{3}$ as $O_{Y}$. Green lines show the valid pair formed by Pauli strings.
IV-C Optimizing Operator Pairing
In the algorithm above, the complexity at the node selection is reduced because we only select two operators $O_{X},O_{Z}$ in each step. The computation complexity is $\binom{N}{2}\sim O(N^{2})$. However, it introduces new overhead traversing to $O_{x}$ and up to $O_{Y}$. In the worst case, the complexity of traversing up and down can be $O(i)$, leading to a total complexity of $O(N^{4})$ (there are $N$ steps).
In this section, we further optimize the algorithm by reducing the traversing time from $O(N)$ to $O(1)$. This improvement is made based on the following statement:
{mdframed}
In step $i$, the $O_{y}$ we find and its ancestor $O_{Y}\in\mathcal{U}$ must always satisfy $O_{y}=desc_{Z}(O_{Y})$
Proof: We first prove two lemmas. {mdframed} Lemma 1: Before step $i$, a leaf $O$ is not paired with another leaf if and only if it is a Z descendant of a node in $\mathcal{U}$. In other words, $\exists O^{\prime}\in\mathcal{U},O=desc_{Z}(O^{\prime})$.
Proof: We prove it by induction.
  1. (I)
    Before step $i=0$, $\mathcal{U}$ includes all leaves. They are all unpaired and Z descendants of themselves.
  2. (II)
    Suppose Lemma 1 is satisfied before step $i$. We prove it holds after step $i$ (holds before step $i+1$).
    After step $i$, we find $O_{X}$, $O_{Y}$, and $O_{Z}$ as the $X,Y$, and $Z$ children of node $O_{2N+1+i}$. Based on the induction hypothesis, all unpaired leaves in the $O_{X},O_{Y}$, and $O_{Z}$ subtree are $O_{x}=desc_{Z}(O_{X}),O_{y}=desc_{Z}(O_{Y})$, and $O_{z}=desc_{Z}(O_{Z})$. Our algorithm pairs $O_{x}$ with $O_{y}$ while leaving $O_{z}$ unpaired. Since $O_{Z}$ is the $Z$ child of $O_{2N+1+i}$, we have $O_{z}=desc_{Z}(O_{Z})=desc_{Z}(O_{2N+1+i}\in\mathcal{U})$ and it is the only unpaired leaf in the $O_{2N+1+i}$ subtree.
    Now, we can check the lemma. $O_{x}$ and $O_{y}$ are paired and no longer Z descendants since their path to $O_{2N+1+i}$ takes $X$ or $Y$ branch. $O_{z}$ is unpaired and still Z descendant of $O_{2N+1+i}\in S$. The lemma holds on $O_{x},O_{y}$, and $O_{z}$. All other leaves are untouched, and the lemma holds on them. $\square$
{mdframed}
Lemma 2: $O_{y}$ is not paired in step $i$.
Proof: Notice that $\forall 0\leq l<N$, leaf $O_{2l}$ must be paired with $O_{2l+1}$. $O_{x}$ is not paired in step $i$ due to Lemma 1. Then, based on how we find $O_{y}$ in Section IV-B, $O_{y}$ must be unpaired. Otherwise, it leads to a contradiction. $\square$ Based on Lemma 1 and $O_{x}=desc_{Z}(O_{X})$, we have $O_{x}$ must be unpaired. Then, based on Lemma 2, $O_{y}$ is also unpaired. Again, by Lemma 1, $O_{y}$ is a Z descendant, indicating $O_{y}=desc_{Z}(O_{Y})$.
The observation hints that the traversing up and down procedure ($\mathtt{desc_{Z}}$ and $\mathtt{traverse\_up}$ in Algorithm 2) only involves the $Z$ descendants and their ancestors. Thus, we could keep two maps: $O\mapsto desc_{Z}(O)$ and $desc_{Z}(O)\mapsto O$ to reduce the complexity of traversing from $O(N)$ to $O(1)$. The map updates when constructing a new parent in each step, as shown in Algorithm 3. Overall, this optimization reduced the total complexity from $O(N^{4})$ to $O(N^{3})$.
Algorithm 3 Optimized Algorithm 2 with $O\mapsto desc_{Z}(O)$ and $desc_{Z}(O)\mapsto O$ maps
[…]
1:$m_{down}[O_{i}]\leftarrow O_{i}$
2:$m_{up}[O_{i}]\leftarrow O_{i}$EndFor
[…]
4$\cdots$ $\triangleright$ Algorithm 2, line $3\sim 19$
5:$O_{2N+1+i}.(X,Y,Z)\leftarrow(O_{X},O_{Y},O_{Z})$
6:$Zdesc\leftarrow m_{down}[O_{Z}]$
7:$m_{down}[O_{2N+1+i}]\leftarrow Zdesc$
8:$m_{up}[Zdesc]\leftarrow O_{2N+1+i}$
9$\cdots$ $\triangleright$ Algorithm 1, line $16$EndFor
10$\cdots$ $\triangleright$ same as Algorithm 1, line $18\sim 20$
11:
[…]
1:return $m_{down}[O]$EndProcedure
2:
[…]
$\mathtt{traverse\_up}$$O,\mathcal{U}$$\triangleright$ $O(1)$ time
1:return $m_{up}[O]$EndProcedure
[…]
TABLE I: Evaluation result of Electronic Structure Model. The best results of each metric are highlighted in bold. ‘–’ indicates the case is too large to solve by Fermihedral (FH). ‘*’ indicates that Fermihedral only finds an approximately optimal solution. ’frz’ indicates freeze core transformation. ’sto3g’ [38] indicates the basis.
Molecule Modes Pauli Weight $CNOT$ Gate Count Circuit Depth
JW BK BTT FH HATT JW BK BTT FH HATT JW BK BTT FH HATT
$H_{2}$ sto3g 4 32 34 36 32 32 21 25 32 22 21 34 25 53 33 34
$LiH$ sto3g frz 6 192 221 225 193* 188 134 211 225 189* 147 191 283 305 254* 198
$LiH$ sto3g 12 3660 3248 3536 3842* 2926 2377 2373 2298 2985* 1642 3174 3249 3306 3987* 2402
$H_{2}O$ sto3g 14 6332 6567 6658 5545 4620 5064 4413 3083 5755 6501 6234 4344
$CH_{4}$ sto3g 18 42476 42646 41530 36983 22798 19051 18645 16304 32951 28704 28326 25745
$O_{2}$ sto3g 20 16904 16828 15364 13076 11653 12126 10382 8912 14532 15426 13914 11331
$NaF$ sto3g 28 247264 218688 207554 192064 102689 93376 89260 69243 141317 141577 133931 110898
$CO_{2}$ sto3g 30 173324 144112 138756 133208 85711 70769 62857 58208 110321 98726 92334 87114
V Evaluation
In this section, we evaluate the proposed Hamiltonian-adaptive Fermion-to-qubit mapping compilation framework HATT against existing Fermion-to-qubit mappings. We also test the scalability and performance of HATT and its optimization.
V-A Benchmark Physics Models
We select the following physical models as our benchmark Fermionic Hamiltonians. They come from various application domains of quantum simulation and have different Fermionic mode coupling structures.
  1. 1.
    Electronic structure model from quantum chmistry [26]. Hamiltonian describes the electrons in a molecule:
    $\mathcal{H}_{\text{e}}=\sum_{p,q}h_{pq}a^{\dagger}_{p}a_{q}+\frac{1}{2}\sum_{p,q,r,s}h_{pqrs}a^{\dagger}_{p}a^{\dagger}_{q}a_{r}a_{s}$
    Geometric data of molecules are from PubChem [23] to determine the coefficients using PySCF [39].
  2. 2.
    Fermi-Hubbard model in condensed-matter physics [1]. Hamiltonian describes a position lattice model of Fermions:
    $\mathcal{H}_{\text{fh}}=\sum_{i,j}\sum_{\sigma=\uparrow,\downarrow}t_{ij}a^{\dagger}_{i,\sigma}a_{j,\sigma}+U\sum_{i}a^{\dagger}_{i,\uparrow}a_{i,\uparrow}a^{\dagger}_{i,\downarrow}a_{i,\downarrow}$
  3. 3.
    Collective neutrino oscillations from astroparticle physics [3, 32, 7]. The Hamiltonian is formulated on a 1D momentum lattice:
    $\begin{split}H_{\nu}=&\sum_{i=1}^{N}{\sum_{a}^{3}{\sqrt{p_{i}^{2}+m_{a}^{2}}}\hat{a}^{\dagger}_{a,i}\hat{a}_{a,i}}\\ +&\sum_{i_{1},i_{2},i_{3}}^{N}\sum_{a,b}^{3}C_{i_{1},i_{2},i_{3}}\hat{a}^{\dagger}_{a,i_{1}}\hat{a}_{a,i_{3}}\hat{a}^{\dagger}_{b,i_{2}}\hat{a}_{b,4}\end{split}$
    where $p_{*,*}$ and $m_{a}$ are the momentum and static mass of neutrino and $C_{i_{1},i_{2},i_{3}}=\mu\left[\hat{p}_{i_{2},x}-\hat{p}_{i_{1},x}\right]\left[\hat{p}_{4,x}-\hat{p}_{i_{3},x}\right]$ ($\mu$ is the two-body coupling strength).
V-B Experiment Setup
V-B1 Implementation
We implemented our Hamiltonian-adaptive Ternary Tree method (HATT) as described in Section III and IV using Python. We leveraged some Pauli operator processing modules from Qiskit [19] and Qiskit Nature [12].
V-B2 Baseline Fermion-to-Qubit Mappings
We compared against a) the Jordan-Wigner transformation (JW[22], and b) the Bravyi-Kitaev transformation (BK[5] in Qiskit Nature [12], c) the balanced ternary tree mapping [27] (BTT), and Fermihedral [25] (FH) that gives optimal Hamiltonian Pauli weight at small scales using a SAT solver.
V-B3 Compilation
The time evolution operator is compiled and optimized with the quantum simulation kernel compiler Paulihedral [24], Rustiq [10] or Tetris [21] followed by Qiskit L3 optimization. We chose $\{CNOT,U3\}$ as the basis gates for noisy simulation. For Tetris [21], we tested against the IBM Manhattan, Sycamore, and Montreal superconducting architectures.
V-B4 Noisy Simulation
We use the Qiskit Aer [19] to simulate circuits with depolarizing noise on single- and two-qubit gates.
V-B5 Real-System Study
We also executed the compiled circuit on the IonQ Forte 1 quantum computer. It has 36 ion trap qubits with all-to-all connectivity, $99.98\%$ single-qubit gate fidelity, $98.99\%$ double-qubit gate fidelity, and $99.02\%$ readout fidelity. Due to current hardware limitations, only the $H_{2}$ molecule simulation is executed.
V-B6 Metrics
We use the following metrics: 1) Pauli weight of the mapped qubit Hamiltonian, 2) $CNOT$ gate count and depth of the compiled circuit, and 3) the simulated system energy in noisy simulation and real-system study.
V-C Pauli Weight and Circuit Metrics
We evaluate the Pauli weight of the qubit Hamiltonians generated by different Fermion-to-qubit mapping methods and the circuits to simulate these Hamiltonians.
V-C1 Electronic Structure Model
Table I shows the Pauli weight and circuit metrics of different molecules. For the smallest case where Fermihedral (FH) can find the optimal Pauli weight, HATT achieves similar results. For larger cases where FH can only provide approximate solutions or even fail to solve, HATT consistently shows the best results in all metrics except for the 6-mode $LiH$ sto3g frz where Jordan-Wigner transformation (JW) is slightly better. Compared with (JW), HATT, on average, reduces Pauli weight by $14.77\%$, number of $CNOT$ gates by $25.84\%$, and circuit depth by $19.33\%$. Compared with the Bravyi-Kitaev transformation (BK), HATT reduces $13.83\%$ Pauli weight, $24.35\%$ $CNOT$ count and $18.91\%$ circuit depth. Compared gainst the balanced ternary tree (BTT), HATT reduces $11.77\%$ Pauli weight, $21.37\%$ $CNOT$ count and $20.58\%$ circuit depth. Compared with JW on more sophisticated workflows, like Rustiq [10] and Tetris [21], HATT still demonstrates better performance for almost all cases (results shown in Table V and IV). Compared with JW+Rustiq, HATT+Rustiq further reduces up to 18.2% $CNOT$ count, 21.83% $U3$ count, and 13.5% circuit depth. For compilation onto superconducting architectures (Manhattan, Sycamore and Montreal), HATT+Tetris can outperform JW+Tetris with reduction on $CNOT$ count, $U3$ count, and circuit depth by up to $17.11\%$, $22.00\%$ and $19.52\%$, respectively. Note that both Rustiq and Tetris are developed using JW. It is possible to further optimize the workflow for the Pauli strings patterns generated by HATT.
TABLE II: Evaluation result of Fermi-Hubbard Model. ‘–’ indicates the case is too large to solve by Fermihedral (FH).
Geometry Modes Pauli Weight $CNOT$ Gate Count Circuit Depth
JW BK BTT FH HATT JW BK BTT FH HATT JW BK BTT FH HATT
$2\times 2$ 8 80 80 86 56 76 51 71 77 37 62 60 99 108 36 73
$2\times 3$ 12 212 200 199 161 187 159 172 161 123 146 161 223 229 160 196
$2\times 4$ 16 304 263 260 230 256 228 183 208 195 189 225 239 256 230 249
$3\times 3$ 18 492 428 408 352 410 378 296 317 266 260 328 391 427 270 356
$2\times 5$ 20 396 348 356 330 287 270 266 224 305 275 320 283
$3\times 4$ 24 704 620 580 524 528 460 433 364 462 529 496 401
$2\times 7$ 28 580 493 502 473 405 380 373 320 465 374 399 333
$3\times 5$ 30 916 756 706 706 661 523 563 490 621 550 629 582
$4\times 4$ 32 1152 790 784 760 842 531 651 483 712 562 749 553
$3\times 6$ 36 1128 932 876 806 794 662 668 544 780 684 694 487
$4\times 5$ 40 1504 1030 986 986 1055 665 782 601 941 561 795 618
TABLE III: Evaluation result of Collective Neutrino Oscillation. Fermihedral (FH) is not included since all the benchmarks in this application are too large for Fermihedral.
Case Modes Pauli Weight $CNOT$ Gate Count Circuit Depth
JW BK BTT HATT JW BK BTT HATT JW BK BTT HATT
$3\times 2F$ 12 1424 1568 1556 1290 776 986 1092 850 1137 1421 1554 1229
$4\times 2F$ 16 4048 4011 4244 3720 2115 2742 2657 2203 3003 3763 3788 3110
$3\times 3F$ 18 5550 5770 5548 5153 2912 3667 3391 2703 3927 4911 4801 3949
$5\times 2F$ 20 9240 9800 9016 7852 4630 5285 5685 4487 6261 7476 8021 6414
$4\times 3F$ 24 16216 16462 14806 14267 7996 8952 8243 7141 10530 12098 11786 10440
$6\times 2F$ 24 18280 18594 16992 15047 8868 9168 9612 8382 11571 13338 13693 11986
$7\times 2F$ 28 32704 31088 28876 25074 15440 14733 15358 13322 19281 21278 22148 19400
$5\times 3F$ 30 37690 33776 32154 31418 17872 17460 16957 15204 21958 23708 23872 21697
$6\times 3F$ 36 75540 66262 60576 58229 34697 30193 29361 26298 41198 41702 41995 38502
$7\times 3F$ 42 136486 114833 101717 99334 60414 48846 48155 45045 69117 67548 67600 64686
V-C2 Fermi-Hubbard Model
Table II shows the Pauli weight and two metrics of the compiled circuits for the Fermi-Hubbard model benchmarks. It can be observed that although FH achieved the minimal Pauli weight for up to $18$ modes, it cannot solve more significant cases. HATT achieves a lower Pauli weight consistently and a lower circuit complexity most of the time against JW, BK, and BTT as our method is tailored according to a specific Hamiltonian and does capture the optimization opportunities within.
On Fermi-Hubbard models, compared to JW, HATT on average reduces Pauli weight by $20.90\%$, number of $CNOT$ gates by $22.90\%$, and circuit depth by $7.88\%$. Compared to BK, HATT reduces $6.48\%$ Pauli weight, $12.11\%$ $CNOT$ count and $8.16\%$ circuit depth. Against BTT, HATT reduces $4.77\%$ Pauli weight, $16.58\%$ $CNOT$ count and $18.11\%$ circuit depth correspondingly. On small scales ($8\sim 18$ modes), HATT shows the results closest to the optimal solution by FH.
V-C3 Collective Neutrino Oscillation
Table III shows the Pauli weight and metrics of compiled circuits of different neutrino oscillation test cases. Fermihedral (FH) is not evaluated since all the cases are too large for FH. Among all the cases, HATT always achieves the lower Pauli weight. Although HATT shows slightly higher circuit overhead compared with JW, the trend shows that HATT has more advantage as the problem size increases. Compared to JW, HATT on average reduces Pauli weight by $15.65\%$, number of $CNOT$ gates by $4.01\%$. Compared to BK, HATT reduces $14.58\%$ Pauli weight, $17.69\%$ $CNOT$ count and $14.92\%$ circuit depth. Against BTT, HATT reduces $11.95\%$ Pauli weight, $17.30\%$ $CNOT$ count and $17.31\%$ circuit depth correspondingly.
TABLE IV: Evaluation result of Electronic Structure Model with Tetris [21] on Manhattan, Sycamore and Montreal
Case $CNOT$ Gate Count $U3$ Gate Count Circuit Depth
JW HATT JW HATT JW HATT
Manhattan
$H_{2}$ sto3g 40 38 47 37 62 56
$H_{2}$ sto3g frz 42 36 40 43 61 57
$H_{2}$ 631g 1479 1420 1302 1172 1744 1659
$H_{2}$ 631g frz 1506 1314 1290 1147 1769 1540
$LiH$ sto3g 7167 6517 5029 4616 7302 6890
$LiH$ sto3g frz 2753 2282 2000 1560 2915 2346
$NH$ sto3g frz 2824 2415 1999 1610 2881 2563
$NH$ 631g 145396 138719 80484 72264 126398 123177
$NH$ 631g frz 85256 80046 48451 46037 76164 73964
$BeH_{2}$ sto3g frz 11646 11513 8911 8980 12711 11959
$O_{2}$ sto3g 48660 46483 24231 23460 38985 39446
Sycamore
$H_{2}$ sto3g 42 38 40 37 60 56
$H_{2}$ sto3g frz 42 38 40 37 56 56
$H_{2}$ 631g 1120 1032 1017 965 1348 1228
$H_{2}$ 631g frz 1122 1040 1044 988 1383 1225
$LiH$ sto3g 5760 5665 4114 3614 6031 5941
$LiH$ sto3g frz 2098 2001 1472 1295 2202 2183
$NH$ sto3g frz 2214 1976 1540 1364 2356 2140
$NH$ 631g 98096 94642 61052 57630 89638 86487
$NH$ 631g frz 60009 56304 38446 33891 56151 52170
$BeH_{2}$ sto3g frz 9607 9350 7155 6895 10427 9805
$O_{2}$ sto3g 31986 29690 19296 17390 28301 27031
Montreal
$H_{2}$ sto3g 42 36 40 43 61 57
$H_{2}$ sto3g frz 42 38 40 37 61 56
$H_{2}$ 631g 1409 1271 1324 1140 1633 1510
$H_{2}$ 631g frz 1435 1417 1300 1160 1699 1641
$LiH$ sto3g 7057 6465 4834 4731 7171 6811
$LiH$ sto3g frz 2865 2522 1949 1628 3013 2553
$NH$ sto3g frz 2862 2467 1939 1524 2999 2523
$NH$ 631g 146957 144169 77300 74026 127955 124624
$NH$ 631g frz 87792 84268 47537 45356 77307 74992
$BeH_{2}$ sto3g frz 11798 12350 8913 8412 12783 12318
$O_{2}$ sto3g 50641 46443 24294 22969 40857 38649
TABLE V: Evaluation result of Electronic Structure Model with Rustiq [10]
Case $CNOT$ Gate Count $U3$ Gate Count Circuit Depth
JW HATT JW HATT JW HATT
$H_{2}$ sto3g 7 7 6 2 7 7
$H_{2}$ sto3g frz 7 7 6 2 7 7
$H_{2}$ 631g 217 227 180 164 164 161
$H_{2}$ 631g frz 217 216 180 170 164 149
$LiH$ sto3g 1348 1230 1031 872 625 561
$LiH$ sto3g frz 423 413 318 306 252 250
$LiH$ 631g 52363 49528 42775 39288 10638 10243
$H_{2}O$ sto3g frz 1914 1793 1580 1481 1926 1748
$H_{2}O$ 631g 114393 111815 109820 106944 96820 95123
$H_{2}O$ 631g frz 107618 104943 86674 81105 20030 19854
$NH$ sto3g 1348 1285 1031 906 625 588
$NH$ sto3g frz 423 346 318 308 252 225
$NH$ 631g frz 22666 18993 17726 13856 5268 4554
$BeH_{2}$ 631g 168863 162917 135486 129317 29096 28040
$BeH_{2}$ 631g frz 106886 102840 82695 79097 20078 19462
$CH_{4}$ sto3g 16641 15741 14429 13343 15074 14140
$CH_{4}$ sto3g frz 10364 10095 7948 7698 3344 3332
$CH_{4}$ 631g 852020 796853 701829 655846 108033 102717
$O_{2}$ sto3g 12012 10592 9526 7529 2766 2582
$O_{2}$ sto3g frz 3155 3103 2313 2374 990 944
$NaF$ sto3g 204673 203415 162845 162837 32126 31928
$NaF$ sto3g frz 6094 5811 5271 5037 5759 5472
V-D Noisy Simulation and Real-System Study
We performed the noisy simulation and real-system study of compiled circuits for $H_{2}$ and $LiH$ sto3g frz cases from the electronic structure model. ‘frz’ means that the inner layer electrons are fixed.
Figure 10: Noisy simulation result of $H_{2}$ and $LiH$ sto3g frz molecule. Energy is measured for 1000 shots. Bias is calculated against the theoretical results.
V-D1 Noisy Simulation
We simulated the circuit generated by different Fermion-to-qubit mappings under depolarizing errors. The error rate range of single-qubit gates is $10^{-5}\sim 10^{-4}$ and $10^{-4}\sim 10^{-3}$ for double-qubit gates. We simulate each circuit for 1000 shots, and the final system energy is measured and compared against the theoretical results.
Figure 10 shows the simulation results of $H_{2}$ (upper half) and $LiH$ sto3g frz (lower half) molecules. Bias and variance are calculated against the theoretical values based on 1000 shots. It can be observed that HATT has the lowest deviation and variance (heatmap closer to blue), which outperforms JW, BK, and BTT, achieving similar results with the optimal FH.
V-D2 Real-System Study
Figure 11 shows the results of $H_{2}$ ground state energy simulation on the IonQ Forte 1 device. The red horizontal lines indicate the average measured energy of 1000 shots using the corresponding mapping. The associated black vertical lines indicate the variance of the measured energy across all shots. The blue line indicates the theoretical result. HATT achieved the second closest to theory, with the closest being the small-scale optimal FH solution. HATT also has the lowest variance.
Figure 11: $H_{2}$ molecule simulation results on IonQ Forte 1 quantum computer
V-E Execution Time and Scalability
To understand the scalability of the proposed tree construction algorithm and the impact of our performance optimization, we compared our unoptimized tree construction algorithm in Section III-C (HATT (unopt)), optimized algorithm in Section IV-C (HATT) and the optimal exhaustive search method Fermihedral [25] (FH) on the time consumption to produce a Fermion-to-qubit mapping for a simple Hamiltonian: $\mathcal{H}_{\mathcal{F}}=\sum_{i=0}^{2N-1}M_{i}$ at different sizes. As shown in Figure 12, both HATT (unopt) and HATT scales to large cases in polynomial time, but FH show its exponential complexity. Also, HATT has a shorter execution time compared with unoptimized HATT (unopt) as our optimization reduces the algorithm’s complexity from $O(N^{4})$ to $O(N^{3})$.
Figure 12: Scalability of Fermihedral (FH), unoptimized HATT (HATT (unopt)) and HATT with optimization (HATT). Black dotted lines (- - -) show the regression of complexity orders.
V-F Performance Cost of Optimized HATT
We also compared the Pauli weight in Table VI of HATT (unopt) and HATT to show how optimization could affect performance. The result shows that HATT and HATT (unopt) generate Hamiltonians with $\sim 0.43\%$ difference in Pauli weight on average. Thus, the optimization does not come with a significant performance trade-off.
TABLE VI: Comparison of Pauli weight between HATT (unopt) and HATT (up to 24 modes)
Case HATT (unopt) HATT Case HATT (unopt) HATT
$H_{2}$ sto3g 32 32 $3\times 3$ 404 410
$LiH$ sto3g frz 188 188 $2\times 5$ 338 330
$LiH$ sto3g 2880 2850 $3\times 4$ 558 524
$H_{2}O$ sto3g 5545 5545 $3\times 2F$ 1266 1290
$CH_{4}$ sto3g 37182 37077 $3\times 3F$ 4976 5153
$O_{2}$ sto3g 13082 13370 $4\times 2F$ 3595 3720
$2\times 2$ 82 76 $4\times 3F$ 14330 14267
$2\times 3$ 194 187 $5\times 2F$ 7844 7852
$2\times 4$ 261 256 $6\times 2F$ 15005 15047
VI Related Works
Fermion-to-Qubit Mappings: Previous Fermion-to-qubit mappings includes the Jordan-Wigner transformation [22], Bravyi-Kitaev transformation [5], parity mapping [4] and ternary tree mapping [20, 27]. The Bravyi-Kitaev transformation and ternary tree mapping generate the theoretical minimum Pauli weight per Majorana operator ($O(\log N)$). However, all of them fail to achieve optimal solutions with a specific Hamiltonian since the pattern of the Fermionic system is disregarded in their construction. Recent superfast encoding [35, 6] captures structures of the Fermionic system to produce optimized mappings. However, it is only restricted to systems with local Fermionic modes and cyclic interaction patterns. Fermihedral [25] solves a given Hamiltonian’s theoretical optimum Pauli weight mapping with an SAT problem. However, SAT is NP-complete, and Fermihedral fails to scale to cases larger than $\sim 20$ qubits. In contrast, HATT combines the goal of optimizing Pauli weight and Hamiltonian information with the ternary tree construction process. It works on any Fermionic system and scales to larger cases with polynomial complexity $O(N^{3})$.
Quantum Simulation Compilers: Some recent works successfully identify the computation pattern in quantum simulation to further optimize quantum circuits, including architectural-aware synthesis [41, 24], reordering Pauli strings for gate cancellation [16, 17], and simultaneous diagonalization [42, 8, 9]. These works happen after Fermion-to-qubit mappings and do not rely on a specific Fermion-to-qubit mapping. HATT is compatible with these works.
Quantum Compilers and Optimizations: Modern quantum compilers, such as Qiskit [19], Cirq [11], Braket [2] and $\ket{\text{ket}}$ [36] apply multiple optimization passes to quantum circuits. These optimizations, including gate cancellations [30] and rewriting [37], qubit routing [28, 29], are applied after the Fermion-to-qubit mapping stage where the Hamiltonian simulation is converted to circuits. Thus, they are compatible with HATT.
VII Conclusion
In this work, we presented the Hamiltonian-Adaptive ternary tree (HATT) compilation framework, a novel optimization for compiling Fermion-to-qubit mapping that leverages the high-level structure of the input Fermionic Hamiltonian. HATT efficiently compiles ternary trees to optimize the mapping process, significantly reducing the Pauli weight and circuit complexity compared to existing methods. Our approach retains the vacuum state preservation property and achieves a polynomial ($O(N^{3})$) complexity, making it suitable for large-scale simulations that classical and exhaustive methods struggle to handle. Extensive evaluations demonstrated that HATT not only achieves close-to-optimal mappings for small systems but also outperforms current constructive methods in larger-scale applications.